OpenStudy (anonymous):

The hypotenuse of a right triangle is 15ft long. One leg is 3ft shorter than the other. Find the lengths of the legs. Please someone help me work through this

OpenStudy (amistre64):

l^2 + (l-3)^2 = 15^2

OpenStudy (angela210793):

15^2=x^2+(x-3)^2

OpenStudy (angela210793):

15^2=x^2+x^2-6x+9 15^2=2x^2-6x+9

OpenStudy (anonymous):

how do i do this amistre64

OpenStudy (amistre64):

well ..... we use the pythag thrm which states: hyp^2 = leg^2 + other leg^2 15^2 = l^2 + (l-3)^2 from the problem stated agree?

OpenStudy (anonymous):

i got x=-9 x=-12 so since there r no - then answer should be 9, 12 right

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

225 = l^2 + l^2 -6l +9 ; go for broke... 0 = 2l^2 -6l - 216 agree?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

factor out the common 2 0 = l^2 -3l -108 ; solve for zeros

OpenStudy (amistre64):

what is your favorite way to get the zeros of this quadratic?

OpenStudy (anonymous):

?

OpenStudy (amistre64):

how do we factor this quadratic?

OpenStudy (amistre64):

we can complete the square if you like

OpenStudy (amistre64):

or we can use the quadratic formula..

OpenStudy (anonymous):

we can use 0=x^2-31x-108

OpenStudy (anonymous):

divide both sides by 31

OpenStudy (anonymous):

x=3,-3 no - x=3 x=3?

OpenStudy (amistre64):

let me clean this up for you; l looks like '1' but its a letter L 0 = x^2 -3x -108 ; try that :)

OpenStudy (anonymous):

divide both sides by 3

OpenStudy (anonymous):

x=36 x=33

OpenStudy (anonymous):

or is it still 12, 9

OpenStudy (amistre64):

108+_____ = x^2 -3x +_____ ; complete the square 108 +9/4 = (x-(3/2))^2 ; sqrt side +- sqrt(441/4) = x-(3/2) ; +3/2 x = 3/2 +- 21/2

OpenStudy (amistre64):

x = 24/2 = 12 or x = -18/2 = -9; lets assume its |-9| = 9 try them out does: 15^2 = 12^2 + 9^2 ??

OpenStudy (anonymous):

225=144=81 YES

OpenStudy (amistre64):

then we found our answers :) You just have to be more efficient at the quads ... and that takes alot of practice ;)

OpenStudy (anonymous):

ok so the answers are 12, 9

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

if leg = 12; the other leg = (12-3) = 9

OpenStudy (anonymous):

ty so much